Help me solve this please
w^4 -12w^2-2=0

Answer by stanbon(75887)

Help me solve this please
w^4 -12w^2-2=0
----
w^2 = [12 +- sqrt(144-4*-2)]/2
--------
w^2 = [12 +- sqrt(152)]/2
--------
w^2 = [12 +- 2sqrt(38)]/2
---
w^2 = 6 + sqrt(38) or w^2 = 6-sqrt(38)
---
Solve for "w".
=================
Cheers,
Stan H.